OpenStudy (anonymous):

answer fast plz! will fan and metal** what is 0.135 as a fraction in simplest form

OpenStudy (anonymous):

200/27

OpenStudy (adajiamcneal):

your right @bubblegumpop<3 pero your backwards ..O.135 = 135/1000 = 27/200